Hi! I nned help with just this one problem. If you could help it wuold be greatly appreciated.
The height of the flagpole is 3/4 the height of the school. The difference in their height is 4.5m. What is the height of each?
I am thankful to all tutors out their who help students like me. You are awesome.
If you can help me that would be great.
Thanks.

Answer by fractalier(2101)

Let x be the height of the school.
Thus (3/4)x is the height of the flagpole.
Since their difference is 4.5, we have
x - (3/4)x = 4.5
(1/4)x = 4.5
x = 18 m
and
(3/4)x = 13.5 m